OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store?

OpenStudy (anonymous):

Set up each equation with the information given. 3c+5d=137 | 2c+7d=161 Solve for one of the variable and plug into the other equation. c= (137-5d)/3 Plugging into the second equation you get 2 * [(137-5d)/3] + 7d = 161 Solve for D. Then solve for C with your newly solved value of D.

OpenStudy (anonymous):

Just read the question again, You can just solve for D in terms of C. skipping that last step. So you solve for D. then plug it in and solve for C.

OpenStudy (anonymous):

okay christ, if you really dont understand the question from there I'm uncertain in your mathematical ablitilies..